OpenStudy (anonymous):

How do I solve \[x^{2} \cos ^{2} y -siny = 0\] ? Can somebody guide me through the steps. I know its implicit differentiation. I just need a bit of guidance.

OpenStudy (anonymous):

So start with that product on the left side.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Use the product rule to deal with x^2(cosy)^2

OpenStudy (anonymous):

dont i need to use chain rule before I can apply the product rule?

OpenStudy (anonymous):

Chain rule comes up in the process of applying the product rule when you differentiate the cosine term.

OpenStudy (radar):

robtobey, take a look at that pigpen problem and let me know if I made an error

OpenStudy (anonymous):

im confused!

OpenStudy (anonymous):

Derivative of the first x^2 times the second (Cosy)^2 + x^2 times the derivative of (cosy)^2.... you would use the chain rule to find the derivative of (Cosy)^2

OpenStudy (anonymous):

I love how errbody helps him and no one else.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

so I got: 2y (dy/dx) (cosy) (-siny) (dy/dx) by the chain rule. is that correct?

OpenStudy (anonymous):

my bad.

OpenStudy (anonymous):

\[2x \cos^{2}y+x^{2} 2y (dy/dx) (cosy) (-\sin) (dy/dx)\]

OpenStudy (anonymous):

is that correct? Please help?

OpenStudy (anonymous):

How did you differentiate (cosy)^2?

OpenStudy (anonymous):

2y (dy/dx) (cosy) (-siny) (dy/dx)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

\[x^2 Cos[y]^2 - Sin[y] == 0 \] \[Cos[y]^2 = 1 - Sin[y]^2 \] \[x^2 (1-\text{Sin}[y]{}^{\wedge}2)-\text{Sin}[y]==0 \] \[x^2-\text{Sin}[y]-x^2 \text{Sin}[y]^2=0 \] Let u equal Sin[y] \[-u+x^2-u^2 x^2=0 \] Solve the above for u \[\left\{u\to \frac{-1-\sqrt{1+4 x^4}}{2 x^2}\right\},\left\{u\to \frac{-1+\sqrt{1+4 x^4}}{2 x^2}\right\} \] Replace u with Sin[y] \[\left\{\text{Sin}[y]\to \frac{-1-\sqrt{1+4 x^4}}{2 x^2}\right\},\left\{\text{Sin}[y]\to \frac{-1+\sqrt{1+4 x^4}}{2 x^2}\right\} \] \[y==\text{ArcSin}\left[\frac{-1-\sqrt{1+4 x^4}}{2 x^2}\right],y==\text{ArcSin}\left[\frac{-1+\sqrt{1+4 x^4}}{2 x^2}\right] \] I hope there no mistakes.

OpenStudy (anonymous):

Think of it as (cos(f(x))^2 where y=f(x). That first y dy/dx shouldnt come out So you get 2(cos(f(x))(-sin(f(x))f'(x). I believe